If a committee of 4 workers is to be chosen what is the probability that it must include at least 2 males?
*there are five males and 5 females

Answer by stanbon(75887)

If a committee of 4 workers is to be chosen what is the probability that it must include at least 2 males?
*there are five males and 5 females
-----
# of ways to succeed : 5C2+5C3+5C4 = 10+10+5 = 25
-----
# of random groups of 4: 10C4 = 210
===============
Ans: 25/210 = 5/42
======================
Cheers,
Stan H.